OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. x - 6y = -15 x - 3y = -3 x - 3(6y - 15) = -3 x - 3(-6y - 15) = -3 (6y - 15) - 3y = -3 (-6y - 15) - 3y = -3

hartnn (hartnn):

x - 6y = -15 add 6y on both sides, what u get ?

OpenStudy (anonymous):

is it x - 3(-6y - 15) = -3?

hartnn (hartnn):

huh ? no. please add 6y to both sides of x - 6y = -15 and tell me what u get?

OpenStudy (anonymous):

x = 6y - 15

hartnn (hartnn):

that is correct! now put that in 2nd equation x - 3y = -3, what u get ?

OpenStudy (anonymous):

(6y - 15) - 3y = -3?

hartnn (hartnn):

absolutely correct. did u understand ?

OpenStudy (anonymous):

yes! thank you so much!

hartnn (hartnn):

welcome ^_^